OpenStudy (anonymous):

21. simplify f(x)=3x^2-4x+1/3x-1. Using complete sentences, explain why f(1)=0, f(0)=1, and f(-1)=-2, yet f(1/3) is undefined. Make sure to show your work.

OpenStudy (anonymous):

= (3x - 1)(x - 1) ----------- (3x - 1) = x - 1

OpenStudy (anonymous):

yes i got that im having difficulting in explaining why f(1)=0, f(0)=1, and f(-1)=-2, yet f(1/3) is undefined

OpenStudy (anonymous):

im sure dumbcrow can explain..

OpenStudy (dumbcow):

Because when x=1/3, the denominator of original f(x) = 0 the limit is (1/3) -1 = -2/3, but at that point it is still undefined, like a hole in the graph

OpenStudy (anonymous):

f(0) = -1

OpenStudy (dumbcow):

yeah i noticed that too...f(0) = -1 for practical purposes the function is x-1 and is treated like a linear function mathematically, it is still undefined at x=1/3